Assuming an efficiency of 49.40%, calculate the actual yield of magnesium nitrate formed from 130.8 g of magnesium and excess copper(II) nitrate.

Mg+Cu(NO3)2⟶Mg(NO3)2+Cu
1
Expert's answer
2019-10-02T08:19:52-0400

Moles of "Mg" reacted"=\frac{130.8g}{24\frac{g}{mol}}=5.45mol"

According to the given efficiency, moles of "Mg(NO_3)_2" formed"=0.4940\times5.45=2.6923mol"

Mass of "Mg(NO_3)_2" formed"=2.6923mol\times148\frac{g}{mol}=398.4604g"